OpenStudy (anonymous):

integral tan^5(t)sec^4(t)dt from 0 to pi/4

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi/4}\tan^5(t)\sec^4(t)dt\]

terenzreignz (terenzreignz):

There's a u-substitution just begging to be made.

OpenStudy (anonymous):

I tried that with tan^5t but then I dont get secant alone

terenzreignz (terenzreignz):

\[\Large \int\limits_{0}^{\frac \pi 4}\tan^5(t)\sec^4(t)dt=\int\limits_{0}^{\frac \pi 4}\tan^5(t)\sec^2(t)\color{red}{\sec^2(t) }dt\]

terenzreignz (terenzreignz):

There's such a thing as 'too ambitious' LOL Try letting u = tan(t) and you'll get results ^_^

OpenStudy (primeralph):